OpenStudy (anonymous):

find the slope -intercept equation for the line with indicated slope and y-intercept ....slope 3/7 , y -intercept (0,-9)

OpenStudy (campbell_st):

same as the last except m = 3/7 and b = -9

OpenStudy (anonymous):

m= 3/7 and c= -9 \[y= mx + c \rightarrow y = \frac{3}{7}x + (-9) \rightarrow y = \frac{3}{7}x-9\] 7y = 3x- 63

OpenStudy (campbell_st):

@dpasingh... the sites for help understanding... not giving answers

OpenStudy (anonymous):

ans is not even correct

OpenStudy (campbell_st):

the y-intercept in always the y value in the point (0, y)

OpenStudy (anonymous):

in slope intercept equation

OpenStudy (anonymous):

@ken_boy05 how you can say that ans is not correct?

OpenStudy (anonymous):

my bad it was simplified i already gave u a medal was looking at just the slope intercept equation before ur simplified form

OpenStudy (anonymous):

dpasignh one more for you